OpenStudy (anonymous):

2y + 11.4 = 2.6 - 0.2y

OpenStudy (anonymous):

Are you supposed to solve for y?

OpenStudy (anonymous):

2y+11.4=2.6-0.2y 2y+0.2y=2.6-11.4 2.2y=-8.8 y=-4 if you are solving for y that ^ is your answer